# Near to Mid-term Risks and Opportunities of Open Source Generative AI

### v1

@@ Abstract
In the next few years, applications of Generative AI are expected to revolutionize a number of different areas, ranging from science & medicine to education. The potential for these seismic changes has triggered a lively debate about potential risks and resulted in calls for tighter regulation, in particular from some of the major tech companies who are leading in AI development. This regulation is likely to put at risk the budding field of open source Generative AI. We argue for the responsible open sourcing of generative AI models in the near and medium term. To set the stage, we first introduce an AI openness taxonomy system and apply it to 40 current large language models. We then outline differential benefits and risks of open versus closed source AI and present potential risk mitigation, ranging from best practices to calls for technical and scientific contributions. We hope that this report will add a much needed missing voice to the current public discourse on near to mid-term AI safety and other societal impact.
@@ 1 Introduction
Parallel to these efforts is a debate on the openness of GenAI models. The digital economy heavily relies on open source software, exemplified by over 60% of global websites using open source servers like Apache and Nginx (Lifshitz-Assaf and Nagle, 2021). This prevalence is underscored by a 2021 European Union report, which concluded that “overall, the [economic] benefits of open source greatly outweigh the costs associated with it” (Blind et al., 2021). Some developers of GenAI models have chosen to openly release trained models (and sometimes data and code too), by leaning on this narrative and claiming that by doing so “[these models] can benefit everyone” and that “it’s safer [to release them]” (Meta, 2023). However, while there has been a flurry of reports and surveys on the impacts of general open source software in areas such as innovation or research within the last few decades (Paulson et al., 2004; Schryen and Kadura, 2009; Von Krogh and Spaeth, 2007), the discourse surrounding the openness of GenAI models presents unique complexities due to the distinctive characteristics of this technology, including e.g., potential dual use and run-away technological progress.
This paper argues that the success of open source in traditional software could be replicated in GenAI with well-defined and followed principles for responsible development and deployment. To this end, we begin by defining different stages of GenAI development/deployment, followed by an empirical analysis of the openness of existing models. With this framework, we then focus on evaluating the risks and opportunities presented by open and closed source GenAI in the near to mid-term. Finally, we make a case for the responsible open sourcing of generative AI models developed in the near to mid-term stages, presenting recommendations to developers on how to achieve this safely and efficiently.
@@ 3 Openness Taxonomy of LLMs
Model developers decide whether to make each component of the training, evaluation and deployment pipeline (Figure 2) private or public, with varying levels of restrictions for the latter. For instance, the developers of LLaMA-2 have publicly released the model architecture and weights, yet they have not shared the code or reward model for Reinforcement Learning from Human Feedback (RLHF) used in the Alignment components (Touvron et al., 2023a). To properly evaluate the openness of each component, we introduce a classification scale for GenAI models in §3.1, which we then apply to 40 high impact LLMs in §3.2. This will help contextualizing the risks and opportunities discussed in §4, and the responsible open sourcing argument we make in §5.
@@ 3.2 Openness Taxonomy of Current LLMs
A full table with the taxonomy of each of the model components is presented in Table LABEL:tab:classification_full (Appendix B). In Figure 4, we show the distribution of openness levels for each of the pipeline components analyzed. Figure 4 clearly shows a balance between open and closed source deployed components (inference code and weights); however, a notable skew exists towards closed source in training data (such as fine-tuning and alignment) and, importantly, in safety evaluation code and data. To fully leverage open source benefits and mitigate risks discussed in the next sections, a significant shift toward responsible development and deployment of open source generative AI is necessary.
@@ 4 Near to Mid-term Risks and Opportunities of Open Source GenAI Models
We describe the risks and opportunities provided by open source models in the near and mid-term (as defined in §2). Our focus is how open source catalyses, minimizes or creates risks and benefits compared to closed source – rather than GenAI in general. Unless stated explicitly, we refer to all artifacts and components of AI when using the term “open source”.
@@ 4.1 Quality and Transparency
* [Uncaptioned image] X Open Models are More Flexible and Customizable Having access to open source models, datasets, and assets significantly aids developers in creating models that are high-performing and specifically tailored to their use-case. Developers have access to far more training approaches, models and datasets. This gives them a powerful starting point when creating a model for a specific application. It also particularly helps cater to less well-resourced languages, domains, and downstream tasks (Bommasani et al., 2023a), as well as enabling personalized models that cater to distinct groups and individuals (Kirk et al., 2023). This has created widespread positive sentiment towards open source, which can be seen in venture capital firm’s significant investment in open-sourcing efforts (Bornstein and Radovanovic, 2023; Horowitz, 2023), and the growing adoption of open source models by companies (Marshall, 2024).
@@ 4.3 Innovation, Industry and Economic Impact
* [Uncaptioned image] X Open Source Empowers Developers and Fosters Innovation Closed source models accessed via an API make product developers reliant on an external provider for essential components of their product or system. This reliance can limit control and maintainability, especially as models can be updated or removed without warning by their owners. Further, with a closed model developers may not own their data or have full control over their data pipeline, which can make it more difficult to innovate on design, steer model performance, change aspects of their system, or understand their own workflows. In contrast, open models offer significant advantages. Developers can modify the model according to their needs, have complete understanding and transparency of the model, and control the data pipeline, which greatly enhances privacy and auditability (Culotta and Mattei, 2023). One important consideration is whether models are released with permissive licenses that suit commercial usecases (see commercial use in §2). This is increasingly common with more recent releases. Open source models could be particularly beneficial in the emerging field of generative AI-powered agents (Chan et al., 2024), where outputs involve performing digital or physical actions (for early examples see Adept’s blog post (AdeptTeam, 2022), and Amazon’s press release (Amazon, 2023)). In this context, product developers are likely to value having more control over models, being able to deploy them on-device, and integrate them in larger, more complex systems.
@@
* [Uncaptioned image] X Open Source Can be More Affordable AI models can enhance individual productivity by automating repetitive and time-consuming tasks, and augmenting workers when completing more complex and high-value tasks. This can help narrow the productivity gap between workers, improving minimum performance standards (Dell’Acqua et al., 2023). In principle, open source AI models increase these benefits as they are available for free. However, substantial operational costs are still involved, such as the staff required to run the models, the time of leadership to organise and oversee their use, and the compute costs for inference (Palazzolo, 2023). Some enterprises might also apply additional protections for security and data to ensure compliance when using open source models, adding further costs. Whether open source is cheaper overall than closed source depends on the maturity and capabilities of the organisation. Generally, larger corporations can bear the overheads involved in open source and overall make substantial savings.
@@
* [Uncaptioned image] X Open Source Can be Easier to Access Open source models are increasingly easy to use and access, with a range of vendors providing SDKs, APIs and downloadable files, such as Replicate, Together, and HuggingFace. Further, they typically require few approvals to start using models, in comparison with more onerous signup processes from closed source providers. One important area where open source lags behind closed source is in providing user interfaces aimed at non-technical audiences. While ChatGPT is easy to interact with and well-known amongst the general public, few open source models have widely-used UIs.
@@ 4.4 Safety
* [Uncaptioned image] X Open Models Can Also be Made to Generate Unsafe Content The flexibility of open source models, as discussed in §4.1, has its drawbacks. Despite their initial alignment, these models can be fine-tuned to produce unsafe content, as exemplified by GPT4Chan and various “uncensored models” on the HuggingFace hub, designed to execute any instruction, irrespective of its safety implications. It is important to recognize, however, that closed models are not impervious to similar risks. Jailbreaks can induce unsafe behaviors in closed models as well (Zou et al., 2023), and recent studies have demonstrated that closed models can easily be fine-tuned to become just as unsafe as open ones (Qi et al., 2023). Nonetheless, ongoing advancements in generative AI safety technology (Dai et al., 2023), particularly through open models, hold the potential for mitigating these risks in the near to mid-term horizon.
@@ 4.5 Societal and Environmental Impact
* [Uncaptioned image] X Open Source Models Can Reduce Energy Use AI model training incurs significant environmental costs from the energy consumption of compute resources. (Strubell et al., 2019; Wu et al., 2022). These impacts, measurable in CO2 emissions, span the entire AI development process, including training and inference.(Verdecchia et al., 2023; Kumar and Davenport, 2023). While accurately quantifying emissions for cloud providers is challenging due to variables like hardware utilization, team practices, geography, and time of day, industry-wide energy consumption can be reduced by sharing of resources that are energy-intensive to create, such as model weights (Saenko, 2023). In addition, open-sourcing can lead to transparent profiling of code to identify energy bottlenecks. This can then be addressed by the community, creating more energy-efficient training methods. For instance, some researchers have put forward small model development paradigms (Schwartz et al., 2019).
@@
* [Uncaptioned image] X Open Source Helps Democratize AI Development Open source empowers developers to utilize resources from major organizations (e.g., companies, governments or research labs), facilitating the reuse of assets and leading to time, effort and money savings. This is crucial for AI development, which is characterized by high costs and complexity, from pre-training models that can cost millions (Knight, 2023) to the creation of expensive human-labeled datasets. This creates a clear societal benefit by enabling non-elites to access and use AI, which can include creating economic opportunities (see §4.3). It is important to acknowledge that, at a higher level, open source models still contain key decisions, datasets and approaches that influence what is built on top of them. In this sense, they are currently undemocratic. They are informed by the values and market priorities of their largely for-profit driven developers.
@@ 5.1 Addressing Common Concerns on Open Sourcing Generative AI
Despite the many benefits of open source, concerns surrounding the increased potential for malicious use, and uncertainty about its societal impact, have prompted calls for keeping generative AI closed source (Seger et al., 2023). There are real risks associated with open source models. However, we believe these are sometimes exaggerated, possibly motivated by the economic interests of market leaders. Most concerns about open sourcing near to mid-term AI models are also pertinent to closed source models.
CLAIM #1: Closed Models Have Inherently Stronger Safeguards than Open Source Models Several studies demonstrate that closed models typically demonstrate fewer safety and security risks, compared to open source (Röttger et al., 2023; Chen et al., 2024; Sun et al., 2024). However, closed models still demonstrate weaknesses, and are particularly vulnerable to jailbreaking techniques (Zou et al., 2023; Chao et al., 2023). Closed model safeguards are easily bypassed through simple manipulations like fine-tuning via accessible services (Qi et al., 2023), prompting the model to repeat a word (Nasr et al., 2023), applying a cypher (Yuan et al., 2023), or instructing the model in another language (Deng et al., 2023; Yong et al., 2023). Completely preventing models from exhibiting undesirable behaviors might not even be possible (Wolf et al., 2023; Petrov et al., 2024). Therefore, it is not clear that closed models are definitively “safer” than open source models. We also anticipate that gaps will narrow over time as open safeguarding methods continue to improve.
@@ 5.2 Recommendations for Safe and Responsible Open Sourcing of Near to Mid-term GenAI Models
To safely and responsibly open source GenAI models, we outline five important priorities for developers, starting with technical recommendations ahead of broader responsibility and socio-technical considerations.
@@
Compare with Closed Source Models Open source models offer advantages like enhanced privacy, customization, transparency, efficiency, and cost-effectiveness. In contrast, commercial closed-source models can stand out in performance, usability, and liability protections. Therefore, comparing the models with their closest commercial closed source alternative is important to quantify, clarify, and understand the trade-offs involved in open sourcing decisions.
@@ 6 Conclusion
The recommendations in §5.2 are a result of combining the openness trends of currently available models in §3.2 with the analysis of §4 on the potential risks and opportunities of open sourcing near to mid-term models. We advocate for the responsible open sourcing of near to mid-term GenAI models. Our work underscores the importance of mitigating risks and addresses prevalent concerns, thereby paving the way for realizing the vast potential benefits open source generative AI offers.
@@ 7 Related Work
The debate around open sourcing GenAI differs from the well-studied impacts of open source software on society (Jaisingh et al., 2008) due to the unique characteristics of the technology. As such, we report related works on two axes: (1) examining the broader impact of GenAI, and (2) on the debate around open sourcing these models.
@@ B.1.1 The EU AI Act
The first comprehensive regulatory framework governing general-purpose AI – including provisions for open-source GenAI – may be the EU AI Act, which is expected to come into full force by 2026 (European Parliament, 2021). The legislation will apply to anyone putting AI services, or their outputs, on the EU market. It guards providers of open-source general-purpose models against risks emanating from downstream use by limiting the providers’ responsibilities to a number of transparency obligations. These transparency obligations include the high-level documentation of training data provenance, as well a specification of intended use cases. Entities deploying GenAI deepfakes are required to disclose their AI-generated nature. These requirements will apply to small business owners to a lesser degree. While comprehensive, the EU AI Act will not apply to recreational or research use and will be superseded by the EU member states’ individual national security interests. Open-source GenAI providers may face additional procedures and obligations if their models are classified as general-purpose AI (GPAI) models of systemic risk, an intentionally vaguely defined criterion that will be adapted as technology progresses. Importantly, the EU AI Act, as perhaps the EU’s first transnational legislation, explicitly affirms the economic benefits of open-source AI.

### v2

# Near to Mid-term Risks and Opportunities of Open-Source Generative AI

@@ Abstract
In the next few years, applications of Generative AI are expected to revolutionize a number of different areas, ranging from science & medicine to education. The potential for these seismic changes has triggered a lively debate about potential risks and resulted in calls for tighter regulation, in particular from some of the major tech companies who are leading in AI development. This regulation is likely to put at risk the budding field of open-source Generative AI. We argue for the responsible open sourcing of generative AI models in the near and medium term. To set the stage, we first introduce an AI openness taxonomy system and apply it to 40 current large language models. We then outline differential benefits and risks of open versus closed source AI and present potential risk mitigation, ranging from best practices to calls for technical and scientific contributions. We hope that this report will add a much needed missing voice to the current public discourse on near to mid-term AI safety and other societal impact.
@@ 1 Introduction
Parallel to these efforts is a debate on the openness of Gen AI models. The digital economy heavily relies on open-source software, exemplified by over 60% of global websites using open-source servers like Apache and Nginx (Lifshitz-Assaf and Nagle, 2021). This prevalence is underscored by a 2021 European Union report, which concluded that “overall, the [economic] benefits of open source greatly outweigh the costs associated with it” (Blind et al., 2021). Some developers of Gen AI models have chosen to openly release trained models (and sometimes data and code too), by leaning on this narrative and claiming that by doing so “[these models] can benefit everyone” and that “it’s safer [to release them]” (Meta, 2023). However, while there has been a flurry of reports and surveys on the impacts of general open-source software in areas such as innovation or research within the last few decades (Paulson et al., 2004; Schryen and Kadura, 2009; Von Krogh and Spaeth, 2007), the discourse surrounding the openness of Gen AI models presents unique complexities due to the distinctive characteristics of this technology, including e.g., potential dual use and run-away technological progress.
This paper argues that the success of open source in traditional software could be replicated in Gen AI with well-defined and followed principles for responsible development and deployment. To this end, we begin by defining different stages of Gen AI development/deployment, followed by an empirical analysis of the openness of existing models through a taxonomy. With this framework, we then focus on evaluating the risks and opportunities presented by open and closed source Gen AI in the near to mid-term. Finally, we make a case for the responsible open sourcing of generative AI models developed in the near to mid-term stages, presenting recommendations to developers on how to achieve this safely and efficiently.
@@ 3 Openness Taxonomy of LLMs
Model developers decide whether to make each component of the training, evaluation and deployment pipeline (Figure 2) private or public, with varying levels of restrictions for the latter. For instance, the developers of LLaMA-2 have publicly released the model architecture and weights, yet they have not shared the code or reward model for Reinforcement Learning from Human Feedback (RLHF) used in the Alignment components (Touvron et al., 2023a). To properly evaluate the openness of each component, we introduce a classification scale for Gen AI models in §3.1, which we then apply to 40 high impact LLMs in §3.2. This will help contextualizing the risks and opportunities discussed in §4, and the responsible open sourcing argument we make in §5. An up-to-date version of the taxonomy of LLMs is also available on this link.
@@ 3.2 Openness Taxonomy of Current LLMs
A full table with the taxonomy of each of the model components is presented in Table LABEL:tab:classification_full (Appendix B). In Figure 4, we show the distribution of openness levels for each of the pipeline components analyzed. Figure 4 clearly shows a balance between open and closed source deployed components (inference code and weights); however, a notable skew exists towards closed source in training data (such as fine-tuning and alignment) and, importantly, in safety evaluation code and data. To fully leverage open source benefits and mitigate risks discussed in the next sections, a significant shift toward responsible development and deployment of open-source generative AI is necessary.
@@ 4 Near to Mid-term Risks and Opportunities of Open Source Gen AI Models
We describe the risks and opportunities provided by open-source models in the near and mid-term (as defined in §2). Our focus is how open source catalyses, minimizes or creates risks and benefits compared to closed source – rather than Gen AI in general. Unless stated explicitly, we refer to all artifacts and components of AI when using the term “open source”.
@@ 4.1 Quality and Transparency
* [Uncaptioned image] X Open Models are More Flexible and Customizable Having access to open-source models, datasets, and assets significantly aids developers in creating models that are high-performing and specifically tailored to their use-case. Developers have access to far more training approaches, models and datasets. This gives them a powerful starting point when creating a model for a specific application. It also particularly helps cater to less well-resourced languages, domains, and downstream tasks (Bommasani et al., 2023a), as well as enabling personalized models that cater to distinct groups and individuals (Kirk et al., 2023). This has created widespread positive sentiment towards open source, which can be seen in venture capital firm’s significant investment in open-sourcing efforts (Bornstein and Radovanovic, 2023; Horowitz, 2023), and the growing adoption of open-source models by companies (Marshall, 2024).
@@ 4.3 Innovation, Industry and Economic Impact
* [Uncaptioned image] X Open Source Empowers Developers and Fosters Innovation Closed source models accessed via an API make product developers reliant on an external provider for essential components of their product or system. This reliance can limit control and maintainability, especially as models can be updated or removed without warning by their owners. Further, with a closed model developers may not own their data or have full control over their data pipeline, which can make it more difficult to innovate on design, steer model performance, change aspects of their system, or understand their own workflows. In contrast, open models offer significant advantages. Developers can modify the model according to their needs, have complete understanding and transparency of the model, and control the data pipeline, which greatly enhances privacy and auditability (Culotta and Mattei, 2023). One important consideration is whether models are released with permissive licenses that suit commercial usecases (see commercial use in §2). This is increasingly common with more recent releases. Open-source models could be particularly beneficial in the emerging field of generative AI-powered agents (Chan et al., 2024), where outputs involve performing digital or physical actions (for early examples see Adept’s blog post (AdeptTeam, 2022), and Amazon’s press release (Amazon, 2023)). In this context, product developers are likely to value having more control over models, being able to deploy them on-device, and integrate them in larger, more complex systems.
@@
* [Uncaptioned image] X Open Source Can be More Affordable AI models can enhance individual productivity by automating repetitive and time-consuming tasks, and augmenting workers when completing more complex and high-value tasks. This can help narrow the productivity gap between workers, improving minimum performance standards (Dell’Acqua et al., 2023). In principle, open-source AI models increase these benefits as they are available for free. However, substantial operational costs are still involved, such as the staff required to run the models, the time of leadership to organise and oversee their use, and the compute costs for inference (Palazzolo, 2023). Some enterprises might also apply additional protections for security and data to ensure compliance when using open-source models, adding further costs. Whether open source is cheaper overall than closed source depends on the maturity and capabilities of the organisation. Generally, larger corporations can bear the overheads involved in open source and overall make substantial savings.
@@
* [Uncaptioned image] X Open Source Can be Easier to Access Open-source models are increasingly easy to use and access, with a range of vendors providing SDKs, APIs and downloadable files, such as Replicate, Together, and HuggingFace. Further, they typically require few approvals to start using models, in comparison with more onerous signup processes from closed source providers. One important area where open source lags behind closed source is in providing user interfaces aimed at non-technical audiences. While ChatGPT is easy to interact with and well-known amongst the general public, few open-source models have widely-used UIs.
@@ 4.4 Safety
* [Uncaptioned image] X Open Models Can Also be Made to Generate Unsafe Content The flexibility of open-source models, as discussed in §4.1, has its drawbacks. Despite their initial alignment, these models can be fine-tuned to produce unsafe content, as exemplified by GPT4Chan and various “uncensored models” on the HuggingFace hub, designed to execute any instruction, irrespective of its safety implications. It is important to recognize, however, that closed models are not impervious to similar risks. Jailbreaks can induce unsafe behaviors in closed models as well (Zou et al., 2023), and recent studies have demonstrated that closed models can easily be fine-tuned to become just as unsafe as open ones (Qi et al., 2023). Nonetheless, ongoing advancements in generative AI safety technology (Dai et al., 2023), particularly through open models, hold the potential for mitigating these risks in the near to mid-term horizon.
@@ 4.5 Societal and Environmental Impact
* [Uncaptioned image] X Open Source Models Can Reduce Energy Use AI model training incurs significant environmental costs from the energy consumption of compute resources. (Strubell et al., 2019; Wu et al., 2022). These impacts, measurable in CO2 emissions, span the entire AI development process, including training and inference (Verdecchia et al., 2023; Kumar and Davenport, 2023). While accurately quantifying emissions for cloud providers is challenging due to variables like hardware utilization, team practices, geography, and time of day, industry-wide energy consumption can be reduced by sharing of resources that are energy-intensive to create, such as model weights (Saenko, 2023). In addition, open-sourcing can lead to transparent profiling of code to identify energy bottlenecks. This can then be addressed by the community, creating more energy-efficient training methods. For instance, some researchers have put forward small model development paradigms (Schwartz et al., 2019).
@@
* [Uncaptioned image] X Open Source Helps Democratize AI Development Open source empowers developers to utilize resources from major organizations (e.g., companies, governments or research labs), facilitating the reuse of assets and leading to time, effort and money savings. This is crucial for AI development, which is characterized by high costs and complexity, from pre-training models that can cost millions (Knight, 2023) to the creation of expensive human-labeled datasets. This creates a clear societal benefit by enabling non-elites to access and use AI, which can include creating economic opportunities (see §4.3). It is important to acknowledge that, at a higher level, open-source models still contain key decisions, datasets and approaches that influence what is built on top of them. In this sense, they are currently undemocratic. They are informed by the values and market priorities of their largely for-profit driven developers.
@@ 5.1 Addressing Common Concerns on Open Sourcing Generative AI
Despite the many benefits of open source, concerns surrounding the increased potential for malicious use, and uncertainty about its societal impact, have prompted calls for keeping generative AI closed source (Seger et al., 2023). There are real risks associated with open-source models. However, we believe these are sometimes exaggerated, possibly motivated by the economic interests of market leaders. Most concerns about open sourcing near to mid-term AI models are also pertinent to closed source models.
CLAIM #1: Closed Models Have Inherently Stronger Safeguards than Open-Source Models Several studies demonstrate that closed models typically demonstrate fewer safety and security risks, compared to open source (Röttger et al., 2023; Chen et al., 2024; Sun et al., 2024). However, closed models still demonstrate weaknesses, and are particularly vulnerable to jailbreaking techniques (Zou et al., 2023; Chao et al., 2023). Closed model safeguards are easily bypassed through simple manipulations like fine-tuning via accessible services (Qi et al., 2023), prompting the model to repeat a word (Nasr et al., 2023), applying a cypher (Yuan et al., 2023), or instructing the model in another language (Deng et al., 2023; Yong et al., 2023). Completely preventing models from exhibiting undesirable behaviors might not even be possible (Wolf et al., 2023; Petrov et al., 2024). Therefore, it is not clear that closed models are definitively “safer” than open-source models. We also anticipate that gaps will narrow over time as open safeguarding methods continue to improve.
@@ 5.2 Recommendations for Safe and Responsible Open Sourcing of Near to Mid-term Gen AI Models
To safely and responsibly open-source Gen AI models, we outline five important priorities for developers, starting with technical recommendations ahead of broader responsibility and socio-technical considerations.
@@
Compare with Closed Source Models Open-source models offer advantages like enhanced privacy, customization, transparency, efficiency, and cost-effectiveness. In contrast, commercial closed-source models can stand out in performance, usability, and liability protections. Therefore, comparing the models with their closest commercial closed source alternative is important to quantify, clarify, and understand the trade-offs involved in open sourcing decisions.
@@ 6 Conclusion
The recommendations in §5.2 are a result of combining the openness trends of currently available models in §3.2 with the analysis of §4 on the potential risks and opportunities of open sourcing near to mid-term models. Following this discussion, we advocate for the responsible open sourcing of near to mid-term Gen AI models.
Note that our position is a balanced one. We advocate that developers should be allowed and encouraged to responsibly open-source Gen AI models developed in the near to mid-term stages, in as much as it makes economic sense for them to do so. Building Gen AI models is an expensive process, and we are sensitive to the argument that for-profit companies should be able to reap some of the financial benefits of their investments in building the technology. Any other position on this matter (e.g., forcing companies to open source their models/pipelines) would seriously risk investment and progress in this area.
However, often for-profit entities will claim open source Gen AI is fundamentally unsafe, and will publicly use this to argue against the open sourcing of these models altogether. This discourages other developers from open sourcing, and we believe this is one of the main factors that contributes to the current skew in the landscape presented in the taxonomy of §3.2 (Figure 4). We reject this argument, and argue in §4 and §5 that (1) there are many benefits that can only be achieved through open sourcing, and (2) the risks are often exaggerated by these for-profit entities. By making these impacts explicit and laying out recommendations for the responsible open sourcing of these models, our aim is to encourage developers to improve the notable skew in Figure 4. This does not mean all models will be open-sourced, only that there would be an improved balance. We note that this should always be voluntary rather than imposed, to avoid disrupting the investment in the area.
Our work underscores the importance of mitigating risks and addresses prevalent concerns, thereby paving the way for realizing the vast potential benefits open-source generative AI offers.
@@ 7 Related Work
The debate around open sourcing GenAI differs from the well-studied impacts of open-source software on society (Jaisingh et al., 2008) due to the unique characteristics of the technology. As such, we report related works on two axes: (1) examining the broader impact of GenAI, and (2) on the debate around open sourcing these models.
@@ B.1.1 The EU AI Act
The first comprehensive regulatory framework governing general-purpose AI – including provisions for open-source Gen AI – may be the EU AI Act, which is expected to come into full force by 2026 (European Parliament, 2021). The legislation will apply to anyone putting AI services, or their outputs, on the EU market for professional purposes, while exempting recreational or academic use, as well as matters relevant to national security. It guards providers of open-source general-purpose models against risks emanating from downstream use by limiting the providers’ responsibilities to a number of transparency obligations. These transparency obligations include the high-level documentation of training data provenance, as well a specification of intended use cases. Entities deploying Gen AI deepfakes are required to disclose their AI-generated nature. These requirements will apply to small business owners to a lesser degree. While comprehensive, the EU AI Act will not apply to recreational or research use and will be superseded by the EU member states’ individual national security interests. Open-source Gen AI providers may face additional procedures and obligations if their models are classified as general-purpose AI (GPAI) models of systemic risk, an intentionally vaguely defined criterion that will be adapted as technology progresses. Importantly, the EU AI Act, as perhaps the EU’s first transnational legislation, explicitly affirms the economic benefits of open-source AI.